I claim that my soda machine dispenses on average more than 12 ounces of soda. Test this claim at a significance level of a=0.05, using the following data: n=81, average=12.2 oz and s=2.4. What is the p-value of the Z test? do not round
Answer #1

since population standard deviation is not known we cannot use z test but we can use t test because we have sample standard deviation value,

t= X - HO 12.2 – 12 s/n 2.4/81 = 0.75

p value = p(t>0.75)@df = 81-1=80 = 0.2277